OpenStudy (laddiusmaximus):

@mertsj find the center and radius of the given circle. x^2+y^2-8x+12y-8=0

OpenStudy (anonymous):

aka complete the square (twice)

OpenStudy (cwrw238):

you can do this by converting your equation to the form (x - a)^2 + (y - b)^2 = r^2 where the center is (a,b) and radius = r you do this by a process called completing the square

OpenStudy (laddiusmaximus):

ok so I have (x^2-8x)+(y^2+12y)=8

OpenStudy (cwrw238):

ok now convert x^6 - 8x and y^2 + 12y to the form (x - a)^2 - c

OpenStudy (cwrw238):

and (y- b)^2 - d

OpenStudy (laddiusmaximus):

(x^2-8x+16)+(y^2+12y+36)=8+16+36

OpenStudy (cwrw238):

x^2 - 8x = (x - 4)^2 - 16

OpenStudy (cwrw238):

and y^2 + 12y = (y + 6)^2 - 36

OpenStudy (cwrw238):

so we get (x -4)^2 + (y + 6)^2 = 8+16+36 = 60

OpenStudy (laddiusmaximus):

gotcha

OpenStudy (cwrw238):

good

OpenStudy (cwrw238):

compare the above withe standard equation and you'll get what you want